# On local Galois deformation rings: generalised reductive groups

### v1

@@ 10. Bounds for the dimensions of reductive subgroups
We may assume that $H$ contains $L^{\prime}$ since otherwise (i) holds by Corollary 10.2. By replacing $H$ by $H\cap L^{0}$ and $L$ by $L^{0}$ we may assume that $L$ is connected. Then $L=Z(L)H$. Let $T$ be a maximal torus in $L$. Then $T$ contains $Z(L)$ and hence $L/H\cong T/(T\cap H)$. The action of $T$ on $\operatorname{Lie}U$ is semisimple, multiplicity free and $(\operatorname{Lie}U)^{T}=0$, [11, 13.18]. Thus $\dim L>\dim H$. If $\dim_{\kappa}((\operatorname{Lie}U)^{*})^{H}\geq 2$, then $\dim_{\kappa}((\operatorname{Lie}U)^{*})^{T\cap H}\geq 2$. Let $\alpha,\beta\in\Phi$ be distinct roots such that the dual characters $\alpha^{*}$ and $\beta^{*}$ appear in $((\operatorname{Lie}U)^{*})^{T\cap H}$. We note that both $\alpha$ and $\beta$ are positive with respect to a Borel subgroup of $G^{\prime}$ containing $T$ and $U$. Thus $\alpha\neq-\beta$. Now $\{\alpha,-\alpha\}$ are the roots of the centraliser of $(\operatorname{Ker}\alpha)^{0}$ in $G^{\prime}$, [11, 13.18(4d)]. Thus $\operatorname{Ker}\alpha\neq\operatorname{Ker}\beta$ and hence $\dim T-\dim(\operatorname{Ker}\alpha\cap\operatorname{Ker}\beta)=2$. Thus $\dim T-\dim T\cap H\geq 2$, as $T\cap H$ is contained in $\operatorname{Ker}\alpha\cap\operatorname{Ker}\beta$. Thus part (i) holds. ∎
@@
11. R-Levi subgroups of codimension $2$
@@
We will let $B_{2}$, $T_{2}$ and $U_{2}$ be subgroups of $\operatorname{PGL}_{2}$ consisting of uppertriangular, diagonal and unipotent uppertriangular matrices, respectively. We will identify $T_{2}$ with $\mathbb{G}_{m}$ via the character $\alpha$, which sends $\bigl{(}\begin{smallmatrix}a&0\\ 0&d\end{smallmatrix}\bigr{)}$ to $ad^{-1}$.
@@
There is a generalised reductive group $G_{1}$ such that $G^{0}_{1}$ is semisimple and
$\overline{G}\cong G_{1}\times\operatorname{PGL}_{2},\quad\overline{P}\cong G_{1}\times B_{2},\quad\overline{L}\cong G_{1}\times T_{2},\quad\overline{U}\cong\{1\}\times U_{2}.$
@@
If $\Phi^{+}(L^{0})$ contains all simple roots of $\Phi^{+}(G^{0})$ then $\Phi(L^{0})=\Phi(G^{0})$ and hence $L^{0}=G^{0}$. Thus there is a simple root $\beta\in\Phi^{+}(G^{0})$, such that $\beta\not\in\Phi^{+}(L^{0})$. It follows from (43), (44) and the assumption $\dim G-\dim L=2$ that $\Phi(G^{0})$ is a disjoint union of $\Phi(L^{0})$ and $\{\beta,-\beta\}$. Let $U_{\alpha}$ be the root subgroup of $G^{0}$ corresponding to $\alpha$. If $\alpha\in\Phi(L^{0})$ then $r\alpha\pm s\beta$ is not a root of $G$ for any integers $r,s>0$, it follows from assertion ()($\ast$)( ∗ ) in the proof of [11, 14.5] that the commutator $(U_{\alpha},U_{\pm\beta})$ vanishes.
We have $\overline{G}{}^{0}\cong G^{0}/Z(G^{0})$, and hence $\overline{G}{}^{0}$ is semisimple by [11, Corollary 14.2(a)]. Let $H_{1}$ be the closed subgroup of $\overline{G}{}^{0}$ generated by the subgroups $U_{\alpha}$ for $\alpha\in\Phi(L^{0})$ and let $H_{2}$ be the closed subgroup of $\overline{G}{}^{0}$ generated by $U_{\beta}$ and $U_{-\beta}$. Arguing as in the proof of [11, Proposition 14.2 (2)] we deduce that $H_{1}$ centralises $H_{2}$, $H_{1}$ and $H_{2}$ generate $\overline{G}{}^{0}$ as a group, and $H_{1}\cap H_{2}$ is contained in the centre of $\overline{G}{}^{0}$. Since the centre of $\overline{G}{}^{0}$ is trivial, we deduce that $\overline{G}{}^{0}\cong H_{1}\times H_{2}$. Moreover, by [20, Lemma 1.2.3] there is a central isogeny $H_{2}\rightarrow\operatorname{PGL}_{2}$. The centre of $H_{2}$ lies in the centre of $\overline{G}{}^{0}$ and hence is trivial, so that $H_{2}\cong\operatorname{PGL}_{2}$.
@@
If $\Theta^{\prime}$ is a subgroup of $\Theta$ then there is a unique subgroup of $H$ of $\operatorname{Aut}(G^{0})$, such that $H$ contains $\overline{G}{}^{0}$, and the image of $H$ in $\operatorname{Aut}(G^{0})/\overline{G}{}^{0}$ is equal to $\Theta^{\prime}$, namely the inverse image of $\Theta^{\prime}$. Uniqueness implies that $H\cong\overline{G}{}^{0}\rtimes\Theta^{\prime}$. Let $\Delta$ be the unique subgroup of $\Theta$, such that $\Delta$ maps isomorphically onto $\overline{G}/\overline{G}{}^{0}\subseteq\operatorname{Aut}(G^{0})/\overline{G}{}^{0}$. Then by the previous remark $\overline{G}\cong\overline{G}{}^{0}\rtimes\Delta$. We claim that every $\delta\in\Delta$ maps $\Phi^{+}(L^{0})$ to itself. The claim implies that $\Delta$ fixes $\beta$, and since it fixes the pinning $\Delta$ acts trivially on $H_{2}$. Thus $\overline{G}\cong(H_{1}\rtimes\Delta)\times\operatorname{PGL}_{2}$ and the proposition follows.
To prove the claim we use the assumption that the map $L/L^{0}\rightarrow G/G^{0}$ is an isomorphism. Thus given $\delta\in\Delta$ we may find $g\in L(\kappa)$, $h_{1}\in H_{1}(\kappa)$ and $h_{2}\in H_{2}(\kappa)$, such that for all $u\in U_{\alpha}$ we have $\delta(u)=(h_{1},h_{2})gug^{-1}(h_{1},h_{2})^{-1}$. The image of $gug^{-1}$ in $\overline{G}{}^{0}$ lies in $H_{1}$ as $L$ normalises $H_{1}$. (It follows from the argument in the proof of [11, Proposition 14.10 (3)] that the closed subgroup of $L$ generated by $U_{\alpha}$ for all $\alpha\in\Phi(L^{0})$ is the derived subgroup of $L^{0}$ and thus is normal in $L$.) Thus $\delta(u)$ lies in $H_{1}$, as $h_{2}$ centralises $H_{1}$. Hence $\delta(\alpha)\neq\pm\beta$ and this finishes the proof of the claim. ∎
@@
If $H_{0}$ is contained in a proper R-parabolic of $L$ then the image of $H_{0}$ is contained in a proper R-parabolic of $\overline{L}$. This would imply that $Z(\overline{L})\overline{H}_{0}$ is contained in a proper R-parabolic of $\overline{L}$, which is not possible as Proposition 11.1 implies that $Z(\overline{L})\overline{H}_{0}=\overline{L}$. Lemma 6.2 in [40] implies that $H_{0}$ is reductive and $Z_{L}(H_{0})^{0}=Z(L)^{0}$. Since $H_{0}\subseteq L$ we have $Z(H_{0})\subseteq Z_{L}(H_{0})$ and hence $Z(H_{0})^{0}\subseteq Z(L)^{0}$. If $\dim Z(H_{0})=\dim Z(L)$ then the quotient would be 00-dimensional and connected and hence $Z(H_{0})^{0}=Z(L)^{0}$. Thus it is enough to produce a cocharacter $\lambda:\mathbb{G}_{m}\rightarrow Z(L)$, such that its image is not contained in $H_{0}$.
We have an exact sequence $0\rightarrow Z(G^{0})\rightarrow L^{0}\rightarrow\overline{L}{}^{0}\rightarrow 0$. Since $L^{0}$ is reductive we have a surjection $Z(L^{0})\twoheadrightarrow Z(\overline{L}{}^{0})$. The group of components of $L$, which we denote by $\Delta$, acts on both groups and induces a finite map on invariants $Z(L^{0})^{\Delta}\rightarrow Z(\overline{L}{}^{0})^{\Delta}$. Hence, the quotient is 00-dimensional and hence every cocharacter $\mu:\mathbb{G}_{m}\rightarrow Z(\overline{L}{}^{0})^{\Delta}$ can be lifted to a cocharacter $\tilde{\mu}:\mathbb{G}_{m}\rightarrow Z(L^{0})^{\Delta}$. Let $\mu:\mathbb{G}_{m}\rightarrow\overline{L}=G_{1}\times\mathbb{G}_{m}$ be the cocharacter $t\mapsto(1,t)$. The image of $\mu$ is contained in $Z^{0}(\overline{L})^{\Delta}$. Its lift $\tilde{\mu}$ takes values in $Z(L^{0})^{\Delta}=Z(L)\cap L^{0}$, see the proof of Proposition 2.9. If the image of $\tilde{\mu}$ is contained in $H_{0}$, then the image of $\mu$ would be contained in $\overline{H}_{0}$, which is not the case. ∎
The following Lemma is essentially [8, Lemma 4.2], but made to work for all $F$ and not just $\mathbb{Q}_{2}$.
@@
[…]
@@
If $\rho_{2}(\Gamma_{F})$ is contained in a maximal split torus of $B_{2}$ then there $u_{2}\in U_{2}(\kappa)$ such that $u_{2}\rho_{2}(\Gamma_{F})u^{-1}_{2}$ is contained in $T_{2}(\kappa)$. Since $L=p_{2}^{-1}(T_{2})$ and $p_{2}:U\overset{\cong}{\longrightarrow}U_{2}$, there is $u\in U(\kappa)$ such that $u\rho(\Gamma_{F})u^{-1}$ is contained in $L(\kappa)$. Since $u^{-1}Lu$ is an R-Levi of $G$ we obtain a contradiction to assumption (1).
@@ 12. Dimension of fibres
Let $x\in X^{\mathrm{gen}}_{P,\rho}(\kappa)$ be the point corresponding to $\rho$ and let $T_{x}X^{\mathrm{gen}}_{P,\rho}$ be the tangent space at $x$. There is natural isomorphism
@@
of $\kappa$-vector spaces, where the action of $\Gamma_{F}$ on $(\operatorname{Lie}U)_{\kappa}$ is given via homomorphism $\rho:\Gamma_{F}\rightarrow L(\kappa)$ composed with the adjoint action of $L$ on $\operatorname{Lie}U$.
@@
The first inequality follows from [8, Lemma 2.2] together with Lemma 12.7. (Its proof shows that $x$ lies on every irreducible component of $X_{P,\rho}^{\mathrm{gen}}$.)
@@
where the disjoint union is taken over all the R-parabolics $P$ of $G$ with R-Levi $L$. We note that there are only finitely many such $P$ by Lemma 2.18. It follows from Lemma 12.9 that (48) induces a surjection on $\overline{\kappa}$-points. Since both the source of and the target of (48) are of finite type over $\overline{\kappa}$, [8, Lemma 3.14] implies that
@@
Let $\mu:\mathbb{G}_{m}\rightarrow G$ be a cocharacter such that $M=L_{\mu}$. Since $\dim G_{\overline{\kappa}}>\dim M$, $\mu$ is not trivial. Since $\dim M=1$, $M^{0}\cong\mathbb{G}_{m}$, and thus $\mu:\mathbb{G}_{m}\rightarrow M^{0}$ is surjective. Hence, $M=Z_{G}(M^{0})$, which means that $M$ centralises $M^{0}$ and so $M^{0}\subseteq Z(M)$. Thus $M^{0}=Z(M)^{0}$ and $\dim Z(M)=\dim M=1$.
Since $M^{0}=M\cap G^{0}_{\overline{\kappa}}$, for the last part we may assume that $G$ is connected. Since $\dim M=1$ and Levi subgroups contain a maximal torus, we may conclude that $M$ is a maximal torus in $G$ and is isomorphic to $\mathbb{G}_{m}$. If $\alpha$ and $\beta$ are distinct positive roots with respect to a Borel subgroup containing $M$ then, as explained in the proof of Lemma 10.6, $\operatorname{Ker}\alpha\neq\operatorname{Ker}\beta$. Since $M\cong\mathbb{G}_{m}$ there is only one positive root. Thus $\dim_{\overline{\kappa}}\operatorname{Lie}G_{\overline{\kappa}}=3$, and so $\dim G_{\overline{\kappa}}=3$. ∎
@@
Moreover, if (49) is an equality then $\dim G_{\overline{k}}-\dim L_{\overline{k}}=2$ and $L$ is $\overline{\rho}^{\mathrm{ss}}$-compatible in the sense of Section 13.1 below. Otherwise,
@@
If (49) is an equality then $\dim L_{\overline{k}}=1$. The last part Lemma 12.17 says that $\dim G_{\overline{\kappa}}=3$. So the difference of dimensions is $2$. The compatibility of $L_{k}$ with $\overline{\rho}^{\mathrm{ss}}$ follows 13.1 below, since $\overline{X}^{\mathrm{git}}_{L,\overline{\rho}^{\mathrm{ss}}}$ is non-empty by construction of $L$.

### v2

@@ 10. Bounds for the dimensions of reductive subgroups
We may assume that $H$ contains $L^{\prime}$ since otherwise (i) holds by Corollary 10.2. By replacing $H$ by $H\cap L^{0}$ and $L$ by $L^{0}$ we may assume that $L$ is connected. Then $L=Z(L)H$. Let $T$ be a maximal torus in $L$. Then $T$ contains $Z(L)$ and hence $L/H\cong T/(T\cap H)$. The action of $T$ on $\operatorname{Lie}U$ is semisimple, multiplicity free and $(\operatorname{Lie}U)^{T}=0$, [10, 13.18]. Thus $\dim L>\dim H$. If $\dim_{\kappa}((\operatorname{Lie}U)^{*})^{H}\geq 2$, then $\dim_{\kappa}((\operatorname{Lie}U)^{*})^{T\cap H}\geq 2$. Let $\alpha,\beta\in\Phi$ be distinct roots such that the dual characters $\alpha^{*}$ and $\beta^{*}$ appear in $((\operatorname{Lie}U)^{*})^{T\cap H}$. We note that both $\alpha$ and $\beta$ are positive with respect to a Borel subgroup of $G^{\prime}$ containing $T$ and $U$. Thus $\alpha\neq-\beta$. Now $\{\alpha,-\alpha\}$ are the roots of the centraliser of $(\operatorname{Ker}\alpha)^{0}$ in $G^{\prime}$, [10, 13.18(4d)]. Thus $\operatorname{Ker}\alpha\neq\operatorname{Ker}\beta$ and hence $\dim T-\dim(\operatorname{Ker}\alpha\cap\operatorname{Ker}\beta)=2$. Thus $\dim T-\dim T\cap H\geq 2$, as $T\cap H$ is contained in $\operatorname{Ker}\alpha\cap\operatorname{Ker}\beta$. Thus part (i) holds. ∎
@@
11. R-Levi subgroups of codimension 22
@@
We will let $B_{2}$, $T_{2}$ and $U_{2}$ be subgroups of $\operatorname{PGL}_{2}$ consisting of upper-triangular, diagonal and unipotent upper-triangular matrices, respectively. We will identify $T_{2}$ with $\mathbb{G}_{m}$ via the character $\alpha$, which sends $\bigl(\begin{smallmatrix}a&0\\ 0&d\end{smallmatrix}\bigr)$ to $ad^{-1}$.
@@
There is a generalised reductive group $G_{1}$ such that $G^{0}_{1}$ is semisimple, $\overline{G}\cong G_{1}\times\operatorname{PGL}_{2}$ and this isomorphism induces isomorphisms:
$\overline{P}\cong G_{1}\times B_{2},\quad\overline{L}\cong G_{1}\times T_{2},\quad\overline{U}\cong\{1\}\times U_{2}.$
@@
If $\Phi^{+}(L^{0})$ contains all simple roots of $\Phi^{+}(G^{0})$ then $\Phi(L^{0})=\Phi(G^{0})$ and hence $L^{0}=G^{0}$. Thus there is a simple root $\beta\in\Phi^{+}(G^{0})$, such that $\beta\not\in\Phi^{+}(L^{0})$. It follows from (43), (44) and the assumption $\dim G-\dim L=2$ that $\Phi(G^{0})$ is a disjoint union of $\Phi(L^{0})$ and $\{\beta,-\beta\}$. Let $U_{\alpha}$ be the root subgroup of $G^{0}$ corresponding to $\alpha$. If $\alpha\in\Phi(L^{0})$ then $r\alpha\pm s\beta$ is not a root of $G$ for any integers $r,s>0$, it follows from assertion $(\ast)$ in the proof of [10, 14.5] that the commutator $(U_{\alpha},U_{\pm\beta})$ vanishes.
We have $\overline{G}{}^{0}\cong G^{0}/Z(G^{0})$, and hence $\overline{G}{}^{0}$ is semisimple by [10, Corollary 14.2(a)]. Let $H_{1}$ be the closed subgroup of $\overline{G}{}^{0}$ generated by the subgroups $U_{\alpha}$ for $\alpha\in\Phi(L^{0})$ and let $H_{2}$ be the closed subgroup of $\overline{G}{}^{0}$ generated by $U_{\beta}$ and $U_{-\beta}$. Arguing as in the proof of [10, Proposition 14.2 (2)] we deduce that $H_{1}$ centralises $H_{2}$, $H_{1}$ and $H_{2}$ generate $\overline{G}{}^{0}$ as a group, and $H_{1}\cap H_{2}$ is contained in the centre of $\overline{G}{}^{0}$. Since the centre of $\overline{G}{}^{0}$ is trivial, we deduce that $\overline{G}{}^{0}\cong H_{1}\times H_{2}$. Moreover, by [20, Lemma 1.2.3] there is a central isogeny $H_{2}\rightarrow\operatorname{PGL}_{2}$. The centre of $H_{2}$ lies in the centre of $\overline{G}{}^{0}$ and hence is trivial, so that $H_{2}\cong\operatorname{PGL}_{2}$.
@@
If $\Theta^{\prime}$ is a subgroup of $\Theta$ then there is a unique subgroup $H$ of $\operatorname{Aut}(G^{0})$, such that $H$ contains $\overline{G}{}^{0}$, and the image of $H$ in $\operatorname{Aut}(G^{0})/\overline{G}{}^{0}$ is equal to $\Theta^{\prime}$, namely the inverse image of $\Theta^{\prime}$. Uniqueness implies that $H\cong\overline{G}{}^{0}\rtimes\Theta^{\prime}$. Let $\Delta$ be the unique subgroup of $\Theta$, such that $\Delta$ maps isomorphically onto $\overline{G}/\overline{G}{}^{0}\subseteq\operatorname{Aut}(G^{0})/\overline{G}{}^{0}$. Then by the previous remark $\overline{G}\cong\overline{G}{}^{0}\rtimes\Delta$. We claim that every $\delta\in\Delta$ maps $\Phi^{+}(L^{0})$ to itself. The claim implies that $\Delta$ fixes $\beta$, and since it fixes the pinning $\Delta$ acts trivially on $H_{2}$. Thus $\overline{G}\cong(H_{1}\rtimes\Delta)\times\operatorname{PGL}_{2}$ and the proposition follows.
To prove the claim we use the assumption that the map $L/L^{0}\rightarrow G/G^{0}$ is an isomorphism. Thus given $\delta\in\Delta$ we may find $g\in L(\kappa)$, $h_{1}\in H_{1}(\kappa)$ and $h_{2}\in H_{2}(\kappa)$, such that for all $u\in U_{\alpha}$ we have $\delta(u)=(h_{1},h_{2})gug^{-1}(h_{1},h_{2})^{-1}$. The image of $gug^{-1}$ in $\overline{G}{}^{0}$ lies in $H_{1}$ as $L$ normalises $H_{1}$. (It follows from the argument in the proof of [10, Proposition 14.10 (3)] that the closed subgroup of $L$ generated by $U_{\alpha}$ for all $\alpha\in\Phi(L^{0})$ is the derived subgroup of $L^{0}$ and thus is normal in $L$.) Thus $\delta(u)$ lies in $H_{1}$, as $h_{2}$ centralises $H_{1}$. Hence $\delta(\alpha)\neq\pm\beta$ and this finishes the proof of the claim. ∎
@@
If $H_{0}$ is contained in a proper R-parabolic of $L$ then the image of $H_{0}$ is contained in a proper R-parabolic of $\overline{L}$. This would imply that $Z(\overline{L})\overline{H}_{0}$ is contained in a proper R-parabolic of $\overline{L}$, which is not possible as Proposition 11.1 implies that $Z(\overline{L})\overline{H}_{0}=\overline{L}$. Lemma 6.2 in [42] implies that $H_{0}$ is reductive and $Z_{L}(H_{0})^{0}=Z(L)^{0}$. Since $H_{0}\subseteq L$ we have $Z(H_{0})\subseteq Z_{L}(H_{0})$ and hence $Z(H_{0})^{0}\subseteq Z(L)^{0}$. If $\dim Z(H_{0})=\dim Z(L)$ then the quotient would be 0-dimensional and connected and hence $Z(H_{0})^{0}=Z(L)^{0}$. Thus it is enough to produce a cocharacter $\lambda:\mathbb{G}_{m}\rightarrow Z(L)$, such that its image is not contained in $H_{0}$.
[…]
The following Lemma is essentially [6, Lemma 4.2], but made to work for all $F$ and not just $\mathbb{Q}_{2}$.
@@
Let $\rho:\Gamma_{F}\rightarrow B_{2}(\kappa)$ be a representation and let $\psi:\Gamma_{F}\rightarrow\kappa^{\times}$ be the character $\psi:\Gamma_{F}\overset{\rho}{\longrightarrow}B_{2}(\kappa)\rightarrow B_{2}(\kappa)/U_{2}(\kappa)\overset{\alpha}{\longrightarrow}\mathbb{G}_{m}(\kappa).$ If the following hold:
@@
If $\rho_{2}(\Gamma_{F})$ is contained in a maximal split torus of $B_{2}$ then there exists $u_{2}\in U_{2}(\kappa)$ such that $u_{2}\rho_{2}(\Gamma_{F})u^{-1}_{2}$ is contained in $T_{2}(\kappa)$. Since $L=p_{2}^{-1}(T_{2})$ and $p_{2}:U\overset{\cong}{\longrightarrow}U_{2}$, there is $u\in U(\kappa)$ such that $u\rho(\Gamma_{F})u^{-1}$ is contained in $L(\kappa)$. Since $u^{-1}Lu$ is an R-Levi of $G$ we obtain a contradiction to assumption (1).
@@ 12. Dimension of fibres
Let $x\in X^{\mathrm{gen}}_{P,\rho}(\kappa)$ be the point corresponding to $\rho$ and let $T_{x}X^{\mathrm{gen}}_{P,\rho}$ be the tangent space at $x$. There is a natural isomorphism
@@
of $\kappa$-vector spaces, where the action of $\Gamma_{F}$ on $(\operatorname{Lie}U)_{\kappa}$ is given via the homomorphism $\rho:\Gamma_{F}\rightarrow L(\kappa)$ composed with the adjoint action of $L$ on $\operatorname{Lie}U$.
@@
The first inequality follows from [6, Lemma 2.2] together with Lemma 12.7. (Its proof shows that $x$ lies on every irreducible component of $X_{P,\rho}^{\mathrm{gen}}$.)
@@
where the disjoint union is taken over all the R-parabolics $P$ of $G$ with R-Levi $L$. We note that there are only finitely many such $P$ by Lemma 2.18. It follows from Lemma 12.9 that (48) induces a surjection on $\overline{\kappa}$-points. Since both the source of and the target of (48) are of finite type over $\overline{\kappa}$, [6, Lemma 3.14] implies that
@@
The assumption on $M$ implies that the semisimple rank of $G^{0}_{\overline{\kappa}}$ is $1$ and hence $G^{0}_{\overline{\kappa}}$ is isomorphic to either $\operatorname{SL}_{2}$ or $\operatorname{PGL}_{2}$, [44, Theorem 20.33], which implies the Lemma. ∎
@@
Moreover, if (49) is an equality [^10] By which we mean both inequalities are equalities. then $\dim G_{\overline{k}}-\dim L_{\overline{k}}=2$ and $L$ is $\overline{\rho}^{\mathrm{ss}}$-compatible in the sense of Section 13.1 below. Otherwise,
@@
If (49) is an equality then $\dim L_{\overline{k}}=1$. The last part of Lemma 12.17 says that $\dim G_{\overline{\kappa}}=3$. So the difference of dimensions is $2$. The compatibility of $L_{k}$ with $\overline{\rho}^{\mathrm{ss}}$ follows 13.1 below, since $\overline{X}^{\mathrm{git}}_{L,\overline{\rho}^{\mathrm{ss}}}$ is non-empty by construction of $L$.